# OSINT Clinic: Co-designing AI-Augmented Collaborative OSINT Investigations for Vulnerability Assessment

### v2

@@ 1. Introduction
Cybersecurity clinics, in which students gain valuable hands-on experience while providing free vulnerability assessments to local organizations, offer a potential solution (con, 2024). The cybersecurity clinic model has recently seen exponential growth thanks to support from private charities, government, and industry (Lederman, [n. d.]). One consortium of cyber clinics reports 880 students participating in cybersecurity training and assisting at least 83 public interest organizations across U.S. (srispens, 2024). However, these clinics also face scalability challenges. Training a diverse group of students to conduct high-quality vulnerability assessments is time-consuming and resource-intensive. Even for trained students, a single investigation can take months. Further, some components of vulnerability assessments, such as controls interviews and facility tours, require significant time investments from clients.
@@
In this paper, we propose the concept of an OSINT Clinic to address the scalability challenges of cyber clinics. Open Source Intelligence, or OSINT, refers to investigations that rely entirely on publicly available data. In a cybersecurity context, OSINT tasks include attack surface mapping for identifying all publicly accessible digital assets such as websites, servers, webcams, and social media accounts, as well as their vulnerabilities; brand monitoring for examining review sites and social media for mentions of the businesses, including negative rumors and disinformation campaigns; and breach data discovery for finding leaked credentials on the dark web or other sources (Hayes and Cappa, 2018; Tabatabaei and Wells, 2017; Browne et al., 2024a). While comprehensive vulnerability assessments sometimes include an OSINT component, cyber clinics rarely, if ever, focus on OSINT.
@@
OSINT provides unique advantages for a student team doing vulnerability assessments. Dealing with publicly available data is a lower risk for both the students and the target businesses; this passive, “zero-touch” approach minimizes accidental damage and saves time for business owners. Second, the ethos of the field dictates that the process for finding, verifying, and documenting each piece of evidence must be transparent and repeatable, providing a rich learning environment for students. Despite the growing interest and promise of OSINT investigations (Hayes and Cappa, 2018; Edwards et al., 2017; Martins and Medeiros, 2022; Urban et al., 2020; Riebe et al., 2023), only a few studies have focused on identifying the specific challenges involved in applying these techniques (Riebe et al., 2023). We address this gap by focusing on identifying and addressing the technical and collaborative challenges with OSINT tasks for vulnerability assessments in the clinic.
@@
OSINT investigations for cybersecurity are complex, requiring sifting through vast amounts of unstructured public data, which can be noisy, overwhelming, and time-consuming (Glassman and Kang, 2012; Belghith et al., 2022; Szymoniak and Foks, 2024; Tabatabaei and Wells, 2017). Analysts often face difficulties in data verification due to the unreliable nature of sources and misinformation (Hassan et al., 2018; Belghith et al., 2022). We argue that generative AI can be effective in addressing these challenges based on previous research on its role in narrowing the skill gap in creative and knowledge work (Doshi and Hauser, 2023; Suri et al., 2024; Wan et al., 2024) and examples from OSINT practitioners and trainers like Matt Edmondson and Chris Poulter (Flashpoint, 2024; Poulter, 2023). Following a long line of CSCW research on bridging the ”socio-technical gap” (Ackerman, 2000), our study aims to effectively integrate generative AI into the workflows of the OSINT clinic and help them scale up.
@@
    What are the challenges of scaling vulnerability assessments in OSINT clinics?
@@
    How can generative AI, particularly foundational models like ChatGPT, address the technical and collaboration challenges of such collaborative OSINT investigations?
To explore these questions and develop design ideas for transforming OSINT-based cybersecurity vulnerability assessments, we designed and launched an OSINT clinic with six undergraduate students. We trained the students on OSINT skills with regular practice sessions and workshops to prepare them for real-world investigations. Over the 2023–24 academic year, we conducted a co-design study by adapting the Matchmaking for AI method (Liu et al., 2024) with three phases (Studies 1–3) to support the work of the students. Study 1 identified key challenges across the Intelligence Cycle phases through workshops, revealing performance gaps in student investigations. In Study 2, we utilized generative AI and collaboration platform probes to explore how AI could address challenges across the stages of planning, data collection, processing, analysis, dissemination, and collaboration. We analyzed the user experience and workshopped design ideas to guide the integration of generative AI into these processes. Finally, Study 3 piloted OSINT-based vulnerability assessments with three small businesses, integrating curated AI prompts into real-world investigations. This study demonstrated AI’s practical benefits and also revealed limitations, such as privacy concerns and difficulties in monitoring progress.
@@
    We make a methodological contribution by adapting the Matchmaking for AI co-design process to include learning goals and collaboration challenges.
@@
    We identify new challenges and learning goals as experienced by participants in the OSINT clinic and explored the use of generative AI to enhance vulnerability assessments. The insights can guide the development of practical, collaborative AI-driven solutions.
@@
    We develop design requirements based on outstanding challenges to support future research in streamlining OSINT investigations.
@@ 2. Related Work
2.1. Vulnerability Assessment using OSINT
A cybersecurity vulnerability assessment is a systematic process of identifying, analyzing, and prioritizing potential security weaknesses in an organization. These assessments can have many components, including internal network discovery and vulnerability scanning with tools like Nmap (nma, 2024) and Nessus (nes, 2024); security controls interviews based on frameworks like CIS (cis, 2024); and physical security reviews looking for access control, fire suppression, and water detection (noa, 2020). Passive information-gathering techniques in OSINT achieve a subset of these tasks to assess cybersecurity vulnerabilities without direct interaction with the target. Prior work informs our use of OSINT tools and techniques relevant to vulnerability assessment (Browne et al., 2024a; Hwang et al., 2022; Pastor-Galindo et al., 2020).
Several models of intelligence processes have been proposed for OSINT investigations (Williams and Blum, 2018; Gibson, 2016; Hwang et al., 2022). We follow Hwang et al’s model to guide our investigations due to its focus on continuous iteration and practical applications to cybersecurity (Hwang et al., 2022). Our approach extends this work by building on crowdsourcing methods to scale up OSINT investigations. Venkatagiri et al. applied the OSINT cycle (Williams and Blum, 2018; Belghith et al., 2022) to create capture-the-flag (CTF) competitions that support collaboratively debunking online misinformation (Venkatagiri et al., 2023). Most closely related to our work, Mukhopadhyay et al. developed a crowdsourcing framework to enable a student crowd trained in the discovery and verification of OSINT to augment ongoing investigations of journalists, fact-checkers, and law enforcement officers (Mukhopadhyay et al., 2024). In contrast to these, our work focuses on scaling up OSINT-based vulnerability assessments by investigating the role of technology and co-designing the effective use of generative AI.
$2.$ CSCW Research and Generative AI Opportunities in OSINT
@@ 2.1. Vulnerability Assessment using OSINT
OSINT investigations are inherently complex, requiring the collection and analysis of large volumes of data as part of a sensemaking process to extract meaningful insights (Venkatagiri et al., 2019; Dailey and Starbird, 2015; Alharthi et al., 2021). Collaboration can enhance these efforts by dividing up workloads and integrating diverse perspectives (Fisher et al., 2012). Previous HCI and CSCW research has studied various approaches to collaborative and crowdsourced investigative work, including top-down strategies (Alcaidinho et al., 2017), bottom-up explorations (Huang et al., 2015), and hybrid methods (Venkatagiri et al., 2021). Belghith et al. studied “social OSINT” — i.e., the growing landscape of competitive and cooperative OSINT-themed events — and characterized it as a community of practice with a culture of transparency and collaboration (Belghith et al., 2022). Existing systems like Newstrition (new, 2022) and CAPER (Aliprandi et al., 2014) facilitate expert collaboration in specific contexts like information verification and prevention of organized crime, but there is potential to expand to non-experts through the use of AI tools (Doroudi et al., 2016; Mukhopadhyay et al., 2024).
AI can support OSINT analysts’ efforts by automating parts of the intelligence cycle, such as using machine learning models for data classification and anomaly detection (Brown et al., 2014; João Rafael Gonçalves Evangelista and Napolitano, 2021; Vinayakumar et al., 2018). However, current AI and machine learning tools in OSINT face challenges due to the lack of focus on real-world applications and inadequate integration with existing intelligence tools (Browne et al., 2024b). Analysts often lack the technical background to understand AI outcomes, which hinders trust and usability (De Vries et al., 2015; McClure Haughey et al., 2020). Moreover, AI solutions fail to address key aspects of intelligence work, such as planning and collaboration (Dorton and Hall, 2021; Browne et al., 2024a). Browne et al. speculate that LLMs can support multiple phases of the Intelligence cycle, but no such tools were included in their review (Browne et al., 2024b). We address the need for developing successful tools through effective integration into the analyst’s workflow and bridging the “socio-technical gap” (Ackerman, 2000).
Following the definition by Sarkar et al., we define generative AI as “an end-user tool whose technical implementation includes a generative model based on deep learning” (Sarkar et al., 2023). Generative AI like large language models (LLMs) can perform complex tasks like reasoning and summarization without fine-tuning, making them potentially transformative for cybersecurity (He et al., 2024b; Mayoral-Vilches et al., 2023). Deng et al. developed PentestGPT, demonstrating that while LLMs are effective in tasks like using testing tools, interpreting outputs, and suggesting next steps in penetration testing, they have difficulty maintaining context throughout the entire process (Deng et al., 2024). Generative AI also offers opportunities for technical assistance by enhancing training and knowledge-sharing among investigators (Chen et al., 2023; Suh et al., 2021). Effective training strategies include demonstrations, practice sessions, feedback, and self-evaluation to build OSINT skills (Mukhopadhyay et al., 2024; Gong et al., 2019; Dow et al., 2012).
@@
Research has begun exploring how generative AI, like LLMs, can enhance collaboration by serving as active participants in group settings (He et al., 2024a; Zhang et al., 2023). Shifting from single human-AI interactions to multi-human and AI environments could improve performance in complex scenarios, providing more context-aware collaboration (Muller and Weisz, 2022; Bansal et al., 2021; He et al., 2024a). Generative AI can facilitate dynamic teamwork and decision-making, improving collaborative sensemaking in high-uncertainty situations (Dorton and Hall, 2021; Suri et al., 2024; He et al., 2024a). Based on related work, we posit that generative AI can significantly improve the effectiveness and scalability of OSINT investigations by focusing on practical integration, collaboration, and training.
@@ 2.3. Co-Designing AI for OSINT
Current AI design paradigms often struggle to identify innovative and feasible design concepts that effectively integrate human values. The Matchmaking for AI approach, introduced by Liu et al. (Liu et al., 2024), addresses this by borrowing from the earlier, more general Matchmaking method (Bly and Churchill, 1999; van Dijk and Zimmerman, 2019). It maps AI techniques to user activities using probes, fostering mutual learning between stakeholders and AI researchers. This method extends traditional co-design practices by focusing on practical feasibility and aligning existing AI capabilities with domain-specific tasks. We propose enhancing this model by incorporating real-world OSINT investigations and collaboration, allowing participants to explore AI’s role in cybersecurity risk assessment.
Generative AI presents valuable material for co-design, enabling stakeholders to explore new ways to enhance OSINT investigations (Sarkar et al., 2023; Kulkarni et al., 2023). Large language models (LLMs) like ChatGPT and Gemini have been used as probes to identify areas for improvement and assess their domain-specific applications (Kulkarni et al., 2023). Such generative models can be integrated into OSINT workflows to support planning (Noy and Zhang, 2023; Bhattacharjee et al., 2023), creative problem-solving (He et al., 2024a; Doshi and Hauser, 2023), and provide real-time feedback (Peng et al., 2023; Wan et al., 2024).

### v3

@@ 1. Introduction
Cybersecurity clinics offer a potential solution (con, 2024). They provide essential cybersecurity services to community organizations — including small businesses, nonprofits, and local governments — while giving students real-world experience. Modeled after legal and medical school clinics, these programs are typically housed at colleges and universities under the direction of clinical professors (cli, 2024b). Students from diverse backgrounds and disciplines receive training to offer free vulnerability assessments and cybersecurity assistance to clients who may lack the resources to secure these services otherwise (con, 2024). This model not only offers a valuable skills-based learning environment for students but also strengthens the cybersecurity resilience of communities. The cybersecurity clinic model has recently seen exponential growth thanks to support from private charities, government, and industry (Lederman, [n. d.]).
@@
However, these clinics also face scalability challenges. Training a diverse group of students to conduct high-quality vulnerability assessments is time-consuming and resource-intensive. Even for trained students, a single investigation can take months. Further, some components of vulnerability assessments, such as controls interviews and facility tours, require significant time investments from clients.
@@
In this paper, we propose the concept of an OSINT Clinic, where students can learn and apply OSINT skills in a real-world context to address the scalability challenges of cybersecurity clinics. Open Source Intelligence, or OSINT, refers to investigations that rely entirely on publicly available data. In a cybersecurity context, OSINT tasks include attack surface mapping for identifying all publicly accessible digital assets such as websites, servers, webcams, and social media accounts, as well as their vulnerabilities; brand monitoring for examining review sites and social media for mentions of the businesses, including negative rumors and disinformation campaigns; and breach data discovery for finding leaked credentials on the dark web or other sources (Hayes and Cappa, 2018; Tabatabaei and Wells, 2017; Browne et al., 2024a). While comprehensive vulnerability assessments sometimes include an OSINT component, cyber clinics rarely, if ever, focus on OSINT.
@@
OSINT provides unique advantages for a student team doing vulnerability assessments. Dealing with publicly available data is a lower risk for both the students and the target businesses; this passive, “zero-touch” approach minimizes accidental damage and saves time for business owners. Second, the ethos of the field dictates that the process for finding, verifying, and documenting each piece of evidence must be transparent and repeatable, providing a rich learning environment for students. Despite the growing interest and promise of OSINT investigations (Hayes and Cappa, 2018; Edwards et al., 2017; Martins and Medeiros, 2022; Urban et al., 2020; Riebe et al., 2023), only a few studies have taken a human-centered approach to investigate specific challenges involved in applying these techniques (Riebe et al., 2023). We address this gap by focusing on identifying and addressing the technical and collaborative challenges with OSINT tasks for vulnerability assessments in the clinic.
@@
OSINT investigations for cybersecurity are complex, requiring sifting through vast amounts of unstructured public data, which can be noisy, overwhelming, and time-consuming (Glassman and Kang, 2012; Belghith et al., 2022; Szymoniak and Foks, 2024; Tabatabaei and Wells, 2017). Analysts often face difficulties in data verification due to the unreliable nature of sources and misinformation (Hassan et al., 2018; Belghith et al., 2022). We argue that generative AI can be effective in addressing these challenges based on previous research on its role in narrowing the skill gap in creative and knowledge work (Doshi and Hauser, 2023; Suri et al., 2024; Wan et al., 2024). Examples from OSINT practitioners and trainers like Matt Edmondson and Chris Poulter (Flashpoint, 2024; Poulter, 2023) also back this up. However, reflecting Ackerman’s “socio-technical gap” (Ackerman, 2000), the technical capabilities of AI currently lag behind the nuanced social requirements of trust and confidentiality for sensitive tasks like OSINT. Following a growing thread of HCI and CSCW research, our study aims to effectively integrate generative AI to augment existing workflows — in this case, those of an OSINT clinic — and help them scale up.
@@
    How can we scale up cybersecurity vulnerability assessments using OSINT?
@@
    What challenges arise when performing vulnerability assessments in the OSINT clinic?
  3. (3)
    How can generative AI, particularly large language models like ChatGPT, address the technical and collaboration challenges of OSINT investigations among non-experts?
  4. (4)
    How can generative AI support real-world OSINT investigations for vulnerability assessments?
To explore these questions, we designed and launched an OSINT clinic with six undergraduate students. We trained the students on OSINT skills with regular practice sessions and workshops to prepare them for real-world investigations. Over the 2023–24 academic year, we conducted a longitudinal co-design study by adapting the Matchmaking for AI method (Liu et al., 2024) with three phases (Studies 1–3) to support the work of the students in the clinic. Study 1 identified key challenges across the Intelligence Cycle phases through workshops, revealing performance gaps in student investigations. In Study 2, we utilized generative AI and collaboration platform probes to explore how AI could help novices address the challenges and develop expertise across the stages of planning, data collection, processing, analysis, and dissemination. We analyzed the user experience and workshopped design ideas to show how generative AI can support training and be integrated into workflows. Finally, Study 3 piloted OSINT-based vulnerability assessments with three small businesses, integrating curated AI prompts into real-world investigations. This study demonstrated AI’s practical benefits and also revealed limitations, such as privacy concerns and difficulties in monitoring progress.
@@
    We make a conceptual contribution by developing the OSINT clinic program to scale up cybersecurity vulnerability assessments based on publicly available information.
@@
    We make a methodological contribution by extending the Matchmaking for AI co-design process to include learning goals, collaboration challenges, and real-world deployments.
@@
    We make an empirical contribution by identifying new challenges and learning goals in the OSINT clinic and addressing them through the use of generative AI.
  4. (4)
    We develop design considerations for practical, AI-augmented, and collaborative investigations based on the use of generative AI.
@@ 2. Related Work
Our paper bridges two key areas of recent HCI research: Clinical Computer Security, which defines protocols for addressing targeted digital security threats (Section 2.1), and CSCW research focused on exploring generative AI tools to facilitate sensemaking and collaboration (Section 2.2). By situating our work at this intersection, we highlight how AI-augmented systems can empower novice investigators in navigating complex OSINT tasks, supporting both their learning goals and the effectiveness of real-world investigations. In terms of method, we extend the Matchmaking for AI co-design framework with a longitudinal study design (Section 2.3), focusing on collaboration and real-world validation. We situate the work within the context of cybersecurity vulnerability assessments using OSINT (Section 2.4).
2.1. Clinical Computer Security in HCI
Clinical Computer Security (CCS) (Freed et al., 2019) is an emerging approach in HCI that provides personalized, hands-on support to individuals facing technology-based security threats. CCS methods involve pairing clients with trained technologists who collaboratively investigate and address digital security vulnerabilities. This process is guided by the Understand-Investigate-Advise (UIA) framework, in which consultants conduct interviews to map a client’s digital footprint, investigate potential security threats (e.g., spyware or compromised accounts), and offer tailored mitigation strategies (Havron et al., 2019). CCS draws inspiration from existing support paradigms such as the Citizen Lab (cit, 2024) and Cybersecurity Clinics (cli, 2024b), which focus on assisting victims of nation-state digital attacks. These clinics not only provide critical services but also serve as research sites for developing new security practices. For example, Tseng et al. co-designed data stewardship approaches by collaborating with consultants and clients in CCS settings (Tseng et al., 2024).
The OSINT Clinic builds on CCS research by applying its foundational principles — such as personalized, collaborative investigations and the UIA framework — to cybersecurity vulnerability assessments for small businesses using OSINT. While CCS has primarily focused on intimate partner violence (IPV) contexts (Freed et al., 2019; Havron et al., 2019; Tseng et al., 2022), the OSINT Clinic demonstrates how these principles can be extended to broader cybersecurity challenges. Beyond tackling a new domain, our work explores the HCI and CSCW challenges of leveraging generative AI to support clinical work. We propose a socio-technical system where AI not only assists with technical tasks but also fosters learning, coordination, and leadership among student analysts.
OSINT investigations face numerous challenges that are amplified in the context of clinical work, where student teams analyze public information. Investigators often work with limited resources such as funds, time, and access to specialized software (Kang and Stasko, 2014; Pastor-Galindo et al., 2020). Issues of data quality, quantity, integration, and ethical considerations further complicate the process (Govardhan et al., 2023). Mukhopadhyay et al. found training on technical skills and ethics was essential to get novices to collaborate successfully with OSINT experts on real-world investigations (Mukhopadhyay et al., 2024). However, comprehensive training is difficult to scale. Effective leadership in OSINT is also challenging as it requires navigating legal and ethical boundaries while applying investigative techniques across various domains (Akhgar et al., 2017). In this paper, we explore how generative AI can help scale this clinic model by understanding and addressing these challenges.
2.2. CSCW Research and Generative AI Opportunities in OSINT
@@
OSINT investigations are inherently complex, requiring the collection and analysis of large volumes of data as part of a sensemaking process to extract meaningful insights (Venkatagiri et al., 2019; Dailey and Starbird, 2015; Alharthi et al., 2021). Collaboration can enhance these efforts by dividing up workloads and integrating diverse perspectives (Fisher et al., 2012). Previous HCI and CSCW research has studied various approaches to collaborative and crowdsourced investigative work, including top-down strategies (Alcaidinho et al., 2017), bottom-up explorations (Huang et al., 2015), and hybrid methods (Venkatagiri et al., 2021). Belghith et al. studied “social OSINT” — i.e., the growing landscape of competitive and cooperative OSINT-themed events — and characterized it as a community of practice with a culture of transparency and collaboration (Belghith et al., 2022). This model with overlapping goals of training and delivering real-world value also closely relates to growing HCI research on how trained and well-supported crowds can handle complex investigations (Venkatagiri et al., 2021, 2023; Mukhopadhyay et al., 2024). Existing systems like Newstrition (new, 2022) and CAPER (Aliprandi et al., 2014) facilitate expert collaboration in specific contexts like information verification and prevention of organized crime, but there is potential to expand to non-experts through the use of AI tools (Doroudi et al., 2016; Mukhopadhyay et al., 2024).
Generative AI presents a promising solution for augmenting OSINT investigations, particularly when compared to custom-built AI solutions (Browne et al., 2024a). Current AI and machine learning tools in OSINT face challenges due to the lack of focus on real-world applications and inadequate integration with existing intelligence tools (Yadav et al., 2023; João Rafael Gonçalves Evangelista and Napolitano, 2021; Browne et al., 2024b). Moreover, AI solutions fail to address key aspects of intelligence work, such as planning and collaboration (Dorton and Hall, 2021; Browne et al., 2024a). Following the definition by Sarkar et al., we define generative AI as “an end-user tool whose technical implementation includes a generative model based on deep learning” (Sarkar et al., 2023). Unlike specialized, narrowly focused AI tools, LLMs are general-purpose and capable of performing a range of complex tasks, such as reasoning, summarization, and web searches, without the need for extensive fine-tuning. Their flexibility makes them well-suited for dynamic and high-uncertainty contexts like OSINT investigations. Prior studies, such as PentestGPT, have demonstrated the effectiveness of LLMs in cybersecurity applications, including interpreting tool outputs and suggesting next steps during penetration testing (Deng et al., 2024).
Generative AI also offers advantages in training and skill development for novice OSINT analysts (Chen et al., 2023; Suh et al., 2021). By supporting iterative learning processes through demonstrations, feedback, and self-evaluation (Mukhopadhyay et al., 2024; Dow et al., 2012), LLMs can help bridge knowledge gaps in areas like tool usage, data collection, and verification. Unlike custom AI solutions, which often require specialized technical expertise to design and deploy, generative AI tools are accessible to a broader audience and can be seamlessly integrated into existing workflows.
@@
Research has begun exploring how generative AI, like LLMs, can enhance collaboration by serving as active participants in group settings (He et al., 2024; Zhang et al., 2023). Shifting from single human-AI interactions to multi-human and AI environments could improve performance in complex scenarios, providing more context-aware collaboration (Muller and Weisz, 2022; Bansal et al., 2021; He et al., 2024). Generative AI can facilitate dynamic teamwork and decision-making, improving collaborative sensemaking in high-uncertainty situations (Dorton and Hall, 2021; Suri et al., 2024; He et al., 2024). Based on related work, we posit that generative AI can significantly improve the effectiveness and scalability of OSINT investigations across the dimensions of practical integration, training, and collaboration.
@@ 2.3. Co-Designing AI for OSINT
Current AI design paradigms often struggle to identify innovative and feasible design concepts that effectively integrate human values. The Matchmaking for AI approach, introduced by Liu et al. (Liu et al., 2024), addresses this by borrowing from the earlier, more general Matchmaking method (Bly and Churchill, 1999; van Dijk and Zimmerman, 2019). It maps AI techniques to user activities using probes, fostering mutual learning between stakeholders and AI researchers. This method extends traditional co-design practices by focusing on practical feasibility and aligning existing AI capabilities with domain-specific tasks.
We extend the Matchmaking for AI method by incorporating a longitudinal aspect, allowing us to study the use of generative AI tools over an extended period and across multiple phases of co-design. This enables AI solutions to be practically integrated into workflows and adaptable to evolving team dynamics (Chin Jr et al., 2009; Stahl, 2005). By embedding the method within a real-world context, we validated design ideas through longitudinal observation of participants conducting vulnerability assessments with actual clients.
2.4. Cybersecurity Vulnerability Assessment using OSINT
The study deals with tasks related to cybersecurity vulnerability assessment, which is a systematic process of identifying, analyzing, and prioritizing potential security weaknesses in an organization. These assessments can have many components, including internal network discovery and vulnerability scanning with tools like Nmap (nma, 2024) and Nessus (nes, 2024); security controls interviews based on frameworks like CIS (cis, 2024); and physical security reviews looking for access control, fire suppression, and water detection (noa, 2020). Passive information-gathering techniques in OSINT achieve a subset of these tasks to assess cybersecurity vulnerabilities without direct interaction with the target. Prior work informs our use of OSINT tools and techniques relevant to vulnerability assessment (Browne et al., 2024a; Hwang et al., 2022; Pastor-Galindo et al., 2020).